# Quantum error correction for long chains of trapped ions

### v2

@@ I Introduction
Fault-tolerant quantum computing relies on extensive quantum error correction to correct faults occurring during the computation before they spread to all the qubits. To maximally benefit from quantum error correction, it is crucial to design quantum error correction schemes that are tailored to the specific constraints of the hardware. Surface codes [1, 2] and color codes [3], which are defined by local measurements on a lattice of qubits, naturally fit with superconducting chips, consisting of a square grid of qubits with nearest neighbor connectivity [4, 5, 6, 7, 8, 9]. Quantum error correction schemes have been optimized for several other hardware platforms such as Majorana qubits [10, 11], neutral atoms [12, 13, 14, 15, 16], cat qubits [17, 18, 19], photonic qubits [20, 21, 22, 23] or short ion chains with shuttling [24, 25, 26].
@@
In this work, we investigate the question of designing quantum error correction schemes for a long chain of trapped ions, which is the core module of a trapped ion quantum computer. The potential of long ion chains has been demonstrated experimentally over the past 10 years with 5-ion chains running Shor’s algorithm [27] or Deutsch–Jozsa and Bernstein-Vazirani algorithms [28], a 14-ion GHZ state [29], a quantum chemistry algorithm executed on a 15-ion chain [30], and quantum error correction experiments with chains of 13 to 16 ions [31, 32, 33]. The performance of a 30-qubit quantum computer based on a long chain of ions is analyzed in [34]. Chains with more than 50 ions have been used for analog quantum simulations [35, 36]. The loading of a chain of 155 ions is reported in [37]. Architectures for large-scale quantum computers based on ion chains connected through shuttling or optical interconnects are proposed in [38, 39, 40]. For a review of trapped ion quantum computing see [41].
@@
The major advantages of long ion chains is their large coherence time and high connectivity. Current trapped ion technologies achieve a gate error rate close to $10^{-3}$ for two-qubit gates and $10^{-4}$ for single-qubit gates [31, 34, 42, 43]. Moreover, qubits stored in a long ion chain are fully connected, in the sense that the native gate set includes entangling gates for every pair of qubits [44, 45]. This high connectivity suggests that trapped ions could be a good fit for the implementation of quantum low-density parity-check (LDPC) codes. Recent simulations show that quantum LDPC codes achieve a significantly lower overhead than surface codes while maintaining the same logical error rate [46, 47, 48]. However, these codes require some form of long-range connectivity that is not available in today’s superconducting chips. This makes long ion chains a natural candidate to host quantum LDPC codes, but a detailed analysis is needed to confirm this intuition. There are major differences between the model considered in previous simulations of quantum LDPC codes and the capability of trapped ion qubits. These differences could translate to a large gap between simulation results and a real implementation of these codes.
@@
Quantum error correction is implemented by executing a quantum circuit called the syndrome extraction circuit. We design syndrome extraction circuits respecting the constraints of the ion chain model (Algorithm 1), that can be used for any stabilizer codes [49]. The high connectivity of the ion chain model offers a lot of flexibility in the construction of this circuit. To exploit this freedom, we propose a syndrome extraction tuning protocol (Algorithm 2) that optimizes the syndrome extraction circuit to achieve a favorable tradeoff between logical error rate and qubit overhead.
To establish a baseline for the ion chain model, we consider surface codes, which achieve state-of-the-art performance for superconducting architecture [1, 2], and bivariate bicycle (BB) codes, which achieve state-of-art-performance when long-range gates are available [47]. We simulate the performance of these codes with a syndrome extraction circuit optimized for the ion chain model. Our baseline family includes surface codes with distance $3,5$ and $7$. The number of BB codes with up to 50 data qubits is massive and some of them have poor parameters. Therefore, we do not simulate all of them. Instead, we simulate a BB code with optimal minimum distance $d$ for each achievable pair of $[[n,k]]$ within the BB code family for $n\leq 50$.
Finally, we study a variant of BB codes defined by weight-5 stabilizer generators, in contrast to the weight-6 stabilizer generators of the original BB codes. We refer to these new codes as BB5 codes. To avoid any confusion, in the rest of this paper we use the term BB6 codes for the original BB codes of [47]. We identify two BB5 codes with parameters $[[30,4,5]]$ and $[[48,4,7]]$ that achieve larger code distance than any BB6 code with the same parameters $[[n,k]]$. For comparison BB6 codes only achieve $[[30,4,4]]$ and $[[48,4,6]]$. The BB5 codes we consider belong to the broader families of tensor-product generalized bicycle codes [50] and multivariate bicycle codes proposed in [51]. In [51] the authors previously discovered a code with the same parameters as our $[[30,4,5]]$ BB5 code by searching over a set of codes that is a subset of BB5 codes.
@@
The rest of this paper is organized as follows. Section II introduces the ion chain model. Syndrome extraction circuits for the ion chain model are designed and optimized in Section III. In Section IV, we investigate the performance of surface codes and BB codes under the ion chain model and we design new codes that outperform these codes.
II The ion chain model
@@
In practice, the ions are confined by an electromagnetic field. Each qubit is stored within two energy levels of an ion. A $n$-qubit chain may contain more than $n$ ions if we use additional ions for sympathetic cooling [52]. Remarkably, operations entangling any of the $\binom{n}{2}$ pairs of qubits are available as native gates in the ion chain model. These operations can be implemented using the Mølmer–Sørensen scheme by illuminating the targeted ions using two laser beams [44, 45]. This allows us to operate on a pair of distant qubits by controlling the laser beams without moving the ions. The operation is mediated by the common vibrations of the ions. See Fig. 1 of [34] for an illustration and for more details on the experimental setup.
III Syndrome extraction circuits for the ion chain model
@@
We focus on stabilizer codes which are defined as the common $+1$-eigenspace of a family of commuting Pauli operators called stabilizer generators [49].
@@
The syndrome extraction circuit is the quantum circuit that performs the measurement of the stabilizer generators of the code. It is executed periodically to correct errors. The ion chain model offers a lot of flexibility when designing syndrome extraction circuits, which can be used to reduce the qubit overhead of quantum error correction. For instance, the all-to-all qubit connectivity enables the use of a single ancilla qubit to sequentially measure all stabilizer generators for any code. This reduces the qubit overhead of the surface and BB6 codes by a factor of almost $2$. Note that this is not possible for quantum hardware with local connectivity constraints, where a single ancilla qubit can only measure stabilizers supported on its neighbors. On the other hand, using a single ancilla qubit may result in a long waiting time and a suboptimal logical error rate. To speed up the syndrome extraction and to reduce the logical error rate, one may use multiple ancilla qubits that are measured simultaneously. In this section, we design a family of syndrome extraction circuits with a variable number of ancilla qubits. Then, we propose a protocol to select the number of ancilla qubits that optimizes the tradeoff between the qubit overhead and the logical error rate.
@@
Input: A list of $n$-qubit Pauli operators $S_{0},S_{1}\dots,S_{r-1}$ supported on qubits $10,1,\dots,n-$ 1. An integer $n_{a}\geq 1$.
@@
The ion chain syndrome extraction circuit is described in Algorithm 1. This circuit fits the constraints of the ion chain model. It measures Pauli operators supported on $n$ data qubits indexed by $10,1,\dots,n-$ 1 using $n_{a}$ ancilla qubits indexed by $n,n+1,\dots,n+n_{a}-1$. The ancilla qubits are measured simultaneously. The parallelization of these measurements yields a significant reduction of the total idle time of the qubits because measurements are the slowest operations in the ion chain model. Other gates remain sequential to respect the ion chain model.
To obtain the syndrome extraction circuit of a stabilizer code with $r$ rounds of syndrome extraction, we set the input of Algorithm 1 as the stabilizer generators of the code repeated $r$ times. In numerical simulations, we use $r=d$ rounds of syndrome extraction where $d$ is the minimum distance of the code, which is a standard choice in the literature. Algorithm 1 is used with an input sequence alternating $X$ and $Z$ operators.
To optimize the syndrome extraction circuit, we need to identify a sweet spot between the number of ancilla qubits consumed and the logical error rate achieved. We propose the syndrome extraction tuning protocol described in Algorithm 2. The protocol keeps increasing the number $n_{a}$ of ancilla qubits as long as each iteration reduces the logical error rate by a factor $\gamma$. The optimized number of ancilla qubits $n_{a}$ returned by Algorithm 2 depends on the code, the measurement time $\tau_{m}$, and to a lesser extent on $p$.
@@
IV Improved codes for the ion chain model
As a baseline, we consider surface codes [1, 2], and BB6 codes [47]. These code constructions are reviewed in Appendix A. For our baseline family, we select surface codes with minimum distance $3,5$ and $7$ and we select a BB6 code achieving the largest minimum distance $d$ for each achievable pair $[[n,k]]$ within BB6 code family for $n\leq 50$. We eliminate codes with $d\leq 2$. The parameters obtained are listed in Table 1.
@@ IV Improved codes for the ion chain model
Table 1: Parameters of baseline codes and BB5 codes. The optimized $n_{a}$ is the number of ancilla qubits used for syndrome extraction, selected by Algorithm 2 with parameters $p=5\times 10^{-4}$, $\tau_{m}=30$ and $\gamma=0.9$.
@@
${\bf{H}}_{X}=[A_{1}+A_{2}~{}|~{}A_{3}+A_{4}+A_{5}],$
${\bf{H}}_{Z}=[A_{3}^{T}+A_{4}^{T}+A_{5}^{T}~{}|~{}A_{1}^{T}+A_{2}^{T}]\cdot$
Rows of ${\bf{H}}_{X}$ and ${\bf{H}}_{Z}$ are respectively the indicator vectors of the $X$ and $Z$ stabilizer generators of the codes [53, 54]. Each $A_{i}$ is an $(m\ell)\times(m\ell)$ permutation matrix of the form $Q_{\ell}^{u}\otimes Q_{m}^{v}$ with $u\in\{0,1,\dots,\ell-1\}$ and $v\in\{0,1,\dots,m-1\}$. Each BB5 code is specified by the values of $\ell,m$ and the matrices $A_{1},\dots,A_{5}$. Table 2 presents two instances of BB5 codes, found by exhaustive search, achieving a larger minimum distance than any BB6 codes with the same parameter $[[n,k]]$. The construction of ${\bf{H}}_{X}$ and ${\bf{H}}_{Z}$ from the matrices $A_{i}$ is identical to Prop. 1 in [51], however we allow for matrices $A_{i}=Q_{\ell}^{u}\otimes Q_{m}^{v}$ whereas the previous construction is restricted to $u=v$. Note the construction of our $[[48,4,7]]$ code uses $u\neq v$ (see Table 2).
[…]
@@
Figure 1: Comparison between BB5 codes and baseline codes. Each point on the blue curve is the minimum logical error rate per logical qubit for a BB6 code from the baseline family.
@@
To compare BB5 codes against our baseline codes, we estimate their logical error rate. In this paper, when we talk about the logical error rate, we always mean the logical error rate per syndrome extraction round and per logical qubit. It is defined as $\frac{q_{X}+q_{Z}}{kd}$ where $k$ is the number of logical qubits and $d$ is minimum distance of the code. The value $q_{Z}$ is defined to be the probability that a $Z$ logical measurement returns a non-trivial outcome in the circuit that initializes all data qubits in the $|0\rangle$ state, performs $d$ rounds of syndrome extraction, measures all the data qubits and extracts the logical outcomes using a decoder. The term $q_{X}$ is defined analogously for $X$ logical measurements, by inserting $H$ gates on the data qubits after initialization and before the final measurement. We use the ion chain syndrome extraction circuit of Algorithm 1. The number of ancilla qubits $n_{a}$ is optimized by Algorithm 2 and reported in Table 1.
The logical error rate is estimated with a Monte Carlo simulation using Stim [55], the BP-OSD decoder [56, 57, 58] for BB5 and BB6 codes, and PyMatching 2 [59] for surface codes. We use BP-OSD with parameters max_bp_iters = 10_000, bp_method = "min_sum", osd_order = 5, osd_method = "osd_cs" . For all the simulations in this paper, the relative error bar is smaller than $6\%$.
Fig. 1 shows that the $[[48,4,7]]$ BB5 code outperforms all the BB6 codes of our baseline. While the $[[49,1,7]]$ surface code exhibits a lower logical error rate per logical qubit at low physical error rates, it encodes $4$ times fewer logical qubits than the $[[48,4,7]]$ BB5 code. Including ancilla qubits, the $[[49,1,7]]$ surface code uses a total of $57$ qubits and the $[[48,4,7]]$ BB5 code uses a total of $54$ qubits. Fig. 2 compares BB5 and BB6 codes with identical $[[n,k]]$ parameters. For both the $[[30,4]]$ and $[[48,4]]$ cases, BB5 codes achieve roughly 6×6$\times$6 × smaller logical error rate than BB6 codes at physical error rate $10^{-3}$. The gap becomes even larger for smaller physical error rates.
V Conclusion

### v3

@@ 1 Introduction
Fault-tolerant quantum computing relies on extensive quantum error correction to correct faults occurring during the computation before they spread to all the qubits. To maximally benefit from quantum error correction, it is crucial to design quantum error correction schemes that are tailored to the specific constraints of the hardware. Surface codes [dennis2002topological, fowler2012surface] and color codes [bombin2006topological], which are defined by local measurements on a lattice of qubits, naturally fit with superconducting chips, consisting of a square grid of qubits with nearest neighbor connectivity [chen2022calibrated, krinner2022realizing, zhao2022realization, google2023suppressing, ai2024quantum, lacroix2024scaling]. Quantum error correction schemes have been optimized for several other hardware platforms such as Majorana qubits [chao2020optimization, paetznick2023performance], neutral atoms [bluvstein2024logical, xu2024constant, veroni2024optimized, reichardt2024logical, pecorari2025high], cat qubits [guillaud2019repetition, ruiz2025ldpc, putterman2025hardware], photonic qubits [bartolucci2023fusion, de2024spin, hilaire2024enhanced, walshe2025linear] or short ion chains with shuttling [ryan2021realization, ryan2024high, paetznick2024demonstration].
@@
In this work, we investigate the question of designing quantum error correction schemes for a long chain of trapped ions, which is the core module of a trapped ion quantum computer. The potential of long ion chains has been demonstrated experimentally over the past 10 years with 5-ion chains running Shor’s algorithm [monz2016realization] or Deutsch–Jozsa and Bernstein-Vazirani algorithms [debnath2016demonstration], a 14-ion GHZ state [monz201114], a quantum chemistry algorithm executed on a 15-ion chain [nam2020ground], and quantum error correction experiments with chains of 13 to 16 ions [egan2021fault, egan2021scaling, postler2022demonstration]. The performance of a 30-qubit quantum computer based on a long chain of ions is analyzed in [chen2024benchmarking]. Chains with more than 50 ions have been used for analog quantum simulations [zhang2017observation, kranzl2022controlling]. The loading of a chain of 155 ions is reported in [kamsap2017experimental]. Architectures for large-scale quantum computers based on ion chains connected through shuttling or optical interconnects are proposed in [kielpinski2002architecture, monroe2014large, schwerdt2024scalable]. For a review of trapped ion quantum computing see [bruzewicz2019trapped].
@@
The major advantages of long ion chains is their large coherence time and high connectivity. Current trapped ion technologies achieve a gate error rate close to $10^{-3}$ for two-qubit gates and $10^{-4}$ for single-qubit gates [egan2021fault, chen2024benchmarking, decross2024computational, loschnauer2024scalable]. Moreover, qubits stored in a long ion chain are fully connected, in the sense that the native gate set includes entangling gates for every pair of qubits [sorensen1999quantum, sorensen2000entanglement]. This high connectivity suggests that trapped ions could be a good fit for the implementation of quantum low-density parity-check (LDPC) codes. Recent simulations show that quantum LDPC codes achieve a significantly lower overhead than surface codes while maintaining the same logical error rate [tremblay2022constant, bravyi2024high, berthusen2025toward]. However, these codes require some form of long-range connectivity that is not available in today’s superconducting chips. This makes long ion chains a natural candidate to host quantum LDPC codes, but a detailed analysis is needed to confirm this intuition. There are major differences between the model considered in previous simulations of quantum LDPC codes and the capability of trapped ion qubits. These differences could translate to a large gap between simulation results and a real implementation of these codes. For example, the fitting parameters for surface code’s logical error rate derived from existing circuit models differ significantly from those obtained using our new model, with the latter more accurately reflecting ion-trap architecture performance. See Appendix˜C for a detailed discussion.
@@
Quantum error correction is implemented by executing a quantum circuit called the syndrome extraction circuit. We design syndrome extraction circuits respecting the constraints of the ion chain model (Algorithm˜1), that can be used for any stabilizer codes [gottesman1997stabilizer]. The high connectivity of the ion chain model offers a lot of flexibility in the construction of this circuit. To exploit this freedom, we propose a syndrome extraction tuning protocol (Algorithm˜2) that optimizes the syndrome extraction circuit to achieve a favorable tradeoff between logical error rate and qubit overhead.
To establish a baseline for the ion chain model, we consider surface codes, which achieve state-of-the-art performance for superconducting architecture [dennis2002topological, fowler2012surface], and bivariate bicycle (BB) codes, which achieve state-of-art-performance when long-range gates are available [bravyi2024high]. We simulate the performance of these codes with a syndrome extraction circuit optimized for the ion chain model. Our baseline family includes surface codes with distance $3,5$ and $7$. The number of BB codes with up to 50 data qubits is massive and some of them have poor parameters. Therefore, we do not simulate all of them. Instead, we simulate a BB code with optimal minimum distance $d$ for each achievable pair of $[[n,k]]$ within the BB code family for $n\leq 50$.
Finally, we study a variant of BB codes defined by weight-5 stabilizer generators, in contrast to the weight-6 stabilizer generators of the original BB codes. We refer to these new codes as BB5 codes. To avoid any confusion, in the rest of this paper we use the term BB6 codes for the original BB codes of [bravyi2024high]. We identify two BB5 codes with parameters $[[30,4,5]]$ and $[[48,4,7]]$ that achieve larger code distance than any BB6 code with the same parameters $[[n,k]]$. For comparison BB6 codes only achieve $[[30,4,4]]$ and $[[48,4,6]]$. The BB5 codes we consider belong to the broader families of tensor-product generalized bicycle codes [kovalev2013quantum] and multivariate bicycle codes proposed in [voss2024multivariate]. In [voss2024multivariate] the authors previously discovered a code with the same parameters as our $[[30,4,5]]$ BB5 code by searching over a set of codes that is a subset of BB5 codes.
@@
The rest of this paper is organized as follows. Section˜2 introduces the ion chain model. Syndrome extraction circuits for the ion chain model are designed and optimized in Section˜3. In Section˜4, we investigate the performance of surface codes and BB codes under the ion chain model and we design new codes that outperform these codes.
2 The ion chain model
@@
Empirical data supports our noise model: [postler2024demonstration] reported a two-qubit gate error rate of $0.027$ and a single-qubit gate error rate of $0.0036$. Another paper [chen2024benchmarking] documented single-qubit gate error rates of $2\times 10^{-4}$ and two-qubit gate error rates of $46.4\times 10^{-4}$. In both cases, single-qubit gate errors were approximately an order of magnitude smaller than two-qubit gate errors. This validates our assumption that single-qubit operations incur an error rate ten times lower than two-qubit operations. It was further reported in [postler2024demonstration] that the measurement and initialization error rates are both $0.003$, which is very close to the single-qubit gate error rate $0.0036$. This supports our choice of assigning the same error rate to single-qubit operations, measurements and initializations in our noise model.
@@
In practice, the ions are confined by an electromagnetic field. Each qubit is stored within two energy levels of an ion. A $n$-qubit chain may contain more than $n$ ions if we use additional ions for sympathetic cooling [myatt1997production]. Remarkably, operations entangling any of the $\binom{n}{2}$ pairs of qubits are available as native gates in the ion chain model. These operations can be implemented using the Mølmer–Sørensen scheme by illuminating the targeted ions using two laser beams [sorensen1999quantum, sorensen2000entanglement]. This allows us to operate on a pair of distant qubits by controlling the laser beams without moving the ions. The operation is mediated by the common vibrations of the ions. See Fig. 1 of [chen2024benchmarking] for an illustration and for more details on the experimental setup.
It is worth noting that most aspects of the ion chain model have been individually highlighted in the literature. For example, the significantly slower measurement times in ion-trap systems is well-known [haffner2008quantum, chamberland2018new], sequential unitary operations have been emphasized [postler2022demonstration], and differential noise rates for two-qubit versus single-qubit operations have been documented [postler2022demonstration, pogorelov2025experimental]. However, our proposed model exhibits a critical modeling advantage over existing works. While prior studies provide only single, fixed data points derived from specific experimental conditions, our model proposes a parameterized framework based on a single variable. This allows us to generate a performance curve across a range of error rates. This methodology directly mirrors the widely accepted parallel circuit model (which assumes uniform noise rates for all operations) used for evaluating code performance on superconducting chips [fowler2012surface, tremblay2022constant, bravyi2024high, berthusen2025toward]. Ultimately, our model provides a comprehensive and computationally efficient tool for simulating fault-tolerant performance on ion chain hardware.
In principle, one could construct a multi-variable noise model for a more precise characterization of ion chain hardware. For example, by assigning three independent variables to denote the error rates for single-qubit operations, measurements, and initializations, instead of using a common $p/10$ rate. However, adopting a multi-variable approach would significantly increase the model’s complexity, making it less convenient for the primary goal of code simulation and performance assessment. We believe our model, parameterized by a single variable $p$, achieves a good trade-off between practical convenience and necessary accuracy. This philosophy is exactly analogous to the widely accepted parallel circuit model used for evaluating code performance on superconducting chips. While the error rates for different operations on superconducting hardware are never precisely uniform, that model’s assumption of uniform noise rates is still considered a reasonable approximation and is used extensively [fowler2012surface, tremblay2022constant, bravyi2024high, berthusen2025toward]. Our model adopts the same pragmatic approach for the ion chain hardware, prioritizing comprehensive, efficient simulation over the marginal precision gained by tracking multiple independent parameters.
3 Syndrome extraction circuits for the ion chain model
@@
We focus on stabilizer codes which are defined as the common +1+1-eigenspace of a family of commuting Pauli operators called stabilizer generators [gottesman1997stabilizer].
@@
The syndrome extraction circuit is the quantum circuit that performs the measurement of the stabilizer generators of the code. It is executed periodically to correct errors. The ion chain model offers a lot of flexibility when designing syndrome extraction circuits, which can be used to reduce the qubit overhead of quantum error correction. For instance, the all-to-all qubit connectivity enables the use of a single ancilla qubit to sequentially measure all stabilizer generators for any code. This reduces the total physical qubit count of the surface and BB6 codes by a factor of almost $2$. Note that this is not possible for quantum hardware with local connectivity constraints, where a single ancilla qubit can only measure stabilizers supported on its neighbors. On the other hand, using a single ancilla qubit may result in a long waiting time and a suboptimal logical error rate. To speed up the syndrome extraction and to reduce the logical error rate, one may use multiple ancilla qubits that are measured simultaneously. In this section, we present a family of syndrome extraction circuits with a variable number of ancilla qubits. Note that this syndrome extraction circuit is a standard application of the Hadamard test; see for example Figure 10.13 on page 473 of [nielsen2010quantum]. Then, we propose a protocol to select the number of ancilla qubits that optimizes the tradeoff between the qubit overhead and the logical error rate.
@@
Input: A list of $n$-qubit Pauli operators $S_{0},S_{1}\dots,S_{r-1}$ supported on qubits $0,1,\dots,n-1$. An integer $n_{a}\geq 1$.
@@
The ion chain syndrome extraction circuit is described in Algorithm˜1. This circuit fits the constraints of the ion chain model. It measures Pauli operators supported on $n$ data qubits indexed by $0,1,\dots,n-1$ using $n_{a}$ ancilla qubits indexed by $n,n+1,\dots,n+n_{a}-1$. The ancilla qubits are measured simultaneously. The parallelization of these measurements yields a significant reduction of the total idle time of the qubits because measurements are the slowest operations in the ion chain model. Other gates remain sequential to respect the ion chain model.
To obtain the syndrome extraction circuit of a stabilizer code with $r$ rounds of syndrome extraction, we set the input of Algorithm˜1 as the stabilizer generators of the code repeated $r$ times. In numerical simulations, we use $r=d$ rounds of syndrome extraction where $d$ is the minimum distance of the code, which is a standard choice in the literature [bravyi2024high]. Algorithm˜1 is used with an input sequence alternating $X$ and $Z$ operators.
To optimize the syndrome extraction circuit, we need to identify a sweet spot between the number of ancilla qubits consumed and the logical error rate achieved. We propose the syndrome extraction tuning protocol described in Algorithm˜2. The protocol keeps increasing the number $n_{a}$ of ancilla qubits as long as each iteration reduces the logical error rate by a factor $\gamma$. A smaller value of $\gamma$ makes the while loop’s condition harder to satisfy, causing the loop to exit earlier and thereby yielding a smaller chosen value for $n_{a}$. In Table˜1, we set $\gamma=0.9$ in Algorithm˜2 to obtain the optimized value of $n_{a}$. We also ran Algorithm˜2 for several $\gamma$ values between $0.8$ and $1$, finding that the algorithm’s output is largely unaffected by $\gamma$. Specifically, for $0.8\leq\gamma\leq 1$, the output varies by at most $2$, and often remains consistent across different $\gamma$ values. The optimized number of ancilla qubits $n_{a}$ returned by Algorithm˜2 depends on the code, the measurement time $\tau_{m}$, and to a lesser extent on $p$. The effectiveness of Algorithm˜2 is validated in Appendix˜B.
As mentioned before, two-qubit gates in ion chains can be implemented via the Mølmer–Sørensen mechanism, where interactions are mediated by the ions’ motional modes [sorensen1999quantum, sorensen2000entanglement]. Generally, the Center-of-Mass (COM) mode is utilized. However, as the length of the ion chain increases, the motional modes become denser in frequency, leading to mode crowding and residual coupling to other modes [haffner2008quantum, wu2018noise, monroe2021programmable]. This residual coupling is a critical limitation because it directly increases the error rate of two-qubit gates, thereby constraining the practical length of ion chains. We note that the precise impact of this effect is beyond the scope of this paper and is not included in our simulation. If accounted for, this effect would likely further reduce the optimal number of ancilla qubits, as it increases physical error rates for longer chains.
@@
4 Improved codes for the ion chain model
Figure 1: Comparison between BB5 codes and baseline codes. Each point on the blue curve is the minimum logical error rate per logical qubit for a BB6 code from the baseline family. More specifically, there are 11 BB6 codes in the baseline family, as listed in Table˜1. For each value of physical error rate $p$, we obtain the logical error rate per logical qubit for these 11 BB6 codes through Monte Carlo simulations, and each point on the blue curve is simply the minimum of these 11 values.
As a baseline, we consider surface codes [dennis2002topological, fowler2012surface], and BB6 codes [bravyi2024high]. These code constructions are reviewed in Appendix˜A. For our baseline family, we select surface codes with minimum distance $3,5$ and $7$ and we select a BB6 code achieving the largest minimum distance $d$ for each achievable pair $[[n,k]]$ within BB6 code family for $n\leq 50$. We eliminate codes with $d\leq 2$. The parameters obtained are listed in Table˜1.
@@ 4 Improved codes for the ion chain model
Table 1: Parameters of baseline codes and BB5 codes. The optimized $n_{a}$ is the number of ancilla qubits used for syndrome extraction, selected by Algorithm˜2 with parameters $p=5\times 10^{-4}$, $\tau_{m}=30$ and $\gamma=0.9$.
@@
${\bf{H}}_{X}=[A_{1}+A_{2}\penalty 10000\ |\penalty 10000\ A_{3}+A_{4}+A_{5}],$
${\bf{H}}_{Z}=[A_{3}^{T}+A_{4}^{T}+A_{5}^{T}\penalty 10000\ |\penalty 10000\ A_{1}^{T}+A_{2}^{T}]\cdot$
Rows of ${\bf{H}}_{X}$ and ${\bf{H}}_{Z}$ are respectively the indicator vectors of the $X$ and $Z$ stabilizer generators of the codes [calderbank1996good, steane1996multiple]. Each $A_{i}$ is an $(m\ell)\times(m\ell)$ permutation matrix of the form $Q_{\ell}^{u}\otimes Q_{m}^{v}$ with $u\in\{0,1,\dots,\ell-1\}$ and $v\in\{0,1,\dots,m-1\}$. Each BB5 code is specified by the values of $\ell,m$ and the matrices $A_{1},\dots,A_{5}$. Table˜2 presents two instances of BB5 codes, found by exhaustive search, achieving a larger minimum distance than any BB6 codes with the same parameter $[[n,k]]$. The construction of ${\bf{H}}_{X}$ and ${\bf{H}}_{Z}$ from the matrices $A_{i}$ is identical to Prop. 1 in [voss2024multivariate], however we allow for matrices $A_{i}=Q_{\ell}^{u}\otimes Q_{m}^{v}$ whereas the previous construction is restricted to $u=v$. Note the construction of our $[[48,4,7]]$ code uses $u\neq v$ (see Table˜2). The paper [voss2024multivariate] introduced a $[[30,4,5]]$ code with stabilizer weight 5 that shares the same code parameters as our $[[30,4,5]]$ BB5 code. However, [voss2024multivariate] did not report any code with parameters matching our $[[48,4,7]]$ BB5 code.
$[[n,k,d]]$ $\ell$ $m$ $A_{1}$ $A_{2}$ $A_{3}$ $A_{4}$ $A_{5}$
@@
To compare BB5 codes against our baseline codes, we estimate their logical error rate. In this paper, when we talk about the logical error rate, we always mean the logical error rate per syndrome extraction round and per logical qubit. It is defined as $\frac{q_{X}+q_{Z}}{kd}$ where $k$ is the number of logical qubits and $d$ is minimum distance of the code. The value $q_{Z}$ is defined to be the probability that a $Z$ logical measurement returns a non-trivial outcome in the circuit that initializes all data qubits in the $|0\rangle$ state, performs $d$ rounds of syndrome extraction, measures all the data qubits and extracts the logical outcomes using a decoder. The term $q_{X}$ is defined analogously for $X$ logical measurements, by inserting $H$ gates on the data qubits after initialization and before the final measurement. We use the ion chain syndrome extraction circuit of Algorithm˜1. The number of ancilla qubits $n_{a}$ is optimized by Algorithm˜2 and reported in Table˜1.
The logical error rate is estimated with a Monte Carlo simulation using Stim [gidney2021stim], the BP-OSD decoder [panteleev2021degenerate, roffe_decoding_2020, Roffe_LDPC_Python_tools_2022] for BB5 and BB6 codes, and PyMatching 2 [Higgott2025sparseblossom] for surface codes. We use BP-OSD with parameters max_bp_iters = 10_000, bp_method = "min_sum", osd_order = 5, osd_method = "osd_cs" . For all the simulations in this paper, the relative error bar is smaller than $6\%$.
Fig.˜1 shows that the $[[48,4,7]]$ BB5 code outperforms all the BB6 codes of our baseline. While the $[[49,1,7]]$ surface code exhibits a lower logical error rate per logical qubit at low physical error rates, it encodes $4$ times fewer logical qubits than the $[[48,4,7]]$ BB5 code. Including ancilla qubits, the $[[49,1,7]]$ surface code uses a total of $57$ qubits and the $[[48,4,7]]$ BB5 code uses a total of $54$ qubits. Fig.˜2 compares BB5 and BB6 codes with identical $[[n,k]]$ parameters. For both the $[[30,4]]$ and $[[48,4]]$ cases, BB5 codes achieve roughly $6\times$ smaller logical error rate than BB6 codes at physical error rate $10^{-3}$. The gap becomes even larger for smaller physical error rates.
5 Conclusion